Prove that the general value of theta which satisfies the equation (costheta+isintheta)×(cos2theta+isintheta).... (Cosntheta+isintheta)=1 is 4mπ over n(n+1) where m is any integer
1
Expert's answer
2020-06-25T18:18:41-0400

(cos"\theta" +isin"\theta" )=ei"^\theta"


(cos2"\theta" +isin2"\theta" )=ei2"^\theta"


(cosn"\theta" +isinn"\theta" )=ein"^\theta"


so

(cos"\theta" +isin"\theta" )×(cos2"\theta" +isin2"\theta" ).... (Cosn"\theta" +isinn"\theta" =1

(ei"^\theta")x(ei2"^\theta").....(ein"^\theta")=1

ei"^\theta"(1+2+...+n)=1

ei"^\theta"{n(n+1)/2}=1

now we write ei"^\theta"  in terms of cos and sin

cos[n(n+1)/2]"\theta" + i sin[n(n+1)/2]"\theta" = 1

for value of "\theta"

cos [n(n+1)/2]"\theta" = 1

[n(n+1)/2 ]"\theta" = cos-1

[n(n+1)/2]"\theta" = 2mπ

"\theta" =4mπ/n(n+1